OpenStudy (anonymous):

Write an equation in standard form parallel to y=3x-2 through the point (5,6)

OpenStudy (anonymous):

do you know what the slope of \[y=3x-2\] is?

OpenStudy (anonymous):

I have no idea.. I'm really bad at this :/

OpenStudy (anonymous):

-2?

OpenStudy (anonymous):

\[\large y = \color{red}mx+\color{blue}b\] has slope \(\color{red}m\) and \(y\) intercept \(\color{blue}b\)

OpenStudy (anonymous):

So it would be y=-2x+3?

OpenStudy (anonymous):

@Reaper534 ??

OpenStudy (anonymous):

? we are still working on finding the slope of \[\large y=\color{red}3x-2\]

OpenStudy (anonymous):

is it 3?

OpenStudy (anonymous):

yes it is !

OpenStudy (anonymous):

Yay! ok so i have to make an equattion with the points 5,6 in it

OpenStudy (anonymous):

Write an equation in standard form PARALLEL to y=3x-2 now means find the equation of the line with slope \(\color{red}3\) through \((5,6)\) right, exactly

OpenStudy (anonymous):

Can you tell me how to do that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok so what do I do first

OpenStudy (anonymous):

write \[\large y-\color{green}{y_1}=\color{red}m(x-\color{blue}{x_1})\] with \[\large \color{red}m=\color{red}3,\color{blue}{x_1}=\color{blue}5,\color{green}{y_1}=\color{green}6\]

OpenStudy (anonymous):

let me know what you get

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

type it in and i will tell you there is another step after that

OpenStudy (anonymous):

Well thats all i really needed.. all i need is to write the equation and thats all

OpenStudy (anonymous):

y−y1=m(x−x1) with m=3,x1=5,y1=6

OpenStudy (anonymous):

no actually you need to write it in standard form that is not standard form

OpenStudy (anonymous):

ok can you help me put it in stardard form?

OpenStudy (anonymous):

i will after you make the replacements and tell me what you get

OpenStudy (anonymous):

you wrote y−y1=m(x−x1) with m=3,x1=5,y1=6 which is what i wrote, explaining how to do it you need to make those replacements

OpenStudy (anonymous):

y-6=3(x-5)

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I have an other question

OpenStudy (anonymous):

that is "point slope" form you need "standard " form the next step is to multiply out on the right

OpenStudy (anonymous):

you may have other questions, and i will help if i can, but this one is not done yet

OpenStudy (anonymous):

y-6=3x-15?

OpenStudy (anonymous):

is that it??

OpenStudy (anonymous):

that is the next step, yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now add \(6\) to both sides to get \(y\) by itself

OpenStudy (anonymous):

y=3x-9

OpenStudy (anonymous):

ok good now it is in "slope intercept" form

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do i divide?

OpenStudy (anonymous):

to put it in "standard form" put the \(x\)'s and \(y\)'x on one side of the equal sign, the number on the other

OpenStudy (anonymous):

no no division

OpenStudy (anonymous):

-3x+y=-9?

OpenStudy (anonymous):

\[y=3x-9 \] is "slope intercept" form \[3x-y=9\] is "standard" form

OpenStudy (anonymous):

ohhhh ok

OpenStudy (anonymous):

yours is good too, but mostly the \(x\) term comes first with a positive coefficient so while \[-3x+y=-9\] is one answer, most people write \[3x-y=9\]

OpenStudy (anonymous):

Ohhhh ok :) thanks! I have an other question

OpenStudy (anonymous):

k go ahead and ask what class is this for?

OpenStudy (anonymous):

solve for f(x)=x^2-4x+3 f=-2

OpenStudy (anonymous):

Algebra ll D:

OpenStudy (anonymous):

Im just practicing with some questions in the book

OpenStudy (anonymous):

is it "find \(f(-2)\)" or is it solve \(f(x)=-2\) ?

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

\[\large f(\diamondsuit)=\diamondsuit^2-4\diamondsuit+3\]

OpenStudy (anonymous):

replace \(\diamondsuit\) by \(-2\) use parentheses

OpenStudy (anonymous):

There are diamonds in the way

OpenStudy (anonymous):

OHHHH

OpenStudy (anonymous):

then sell them

OpenStudy (anonymous):

Ok i did that. Do i multiply

OpenStudy (anonymous):

you compute \[(-2)^2-4\times (-2)+3\]

OpenStudy (anonymous):

WHat?

OpenStudy (anonymous):

\[f(x)=x^2-4x+3 \] \[f(-2)=(-2)^2-4\times (-2)+3 \]

OpenStudy (anonymous):

let me know when you get \(15\)